OpenStudy (anonymous):

when you have a term like (-2^-x)ln2 and you want to bring the exponent down to make it positive how come its -(1/(2^x)(ln2) instead of -(ln2/(2^x))?

OpenStudy (anonymous):

why does \[-2^{-x}\ln2\]=\[ -1\over 2^{x}\ln2\] and not \[\ln2 \over 2^{x}\]?

OpenStudy (anonymous):

because \[2^{-x}=\frac{1}{2^x}\]

OpenStudy (anonymous):

should be \[-2^{-x}\ln(x)=-\frac{\ln(x)}{2^x}\]

OpenStudy (anonymous):

I'm with you. ln2 is just a number. So I think it should be \[-\ln 2/2^{x}\] I'm curious where you saw it as \[-1/2^{x}\ln 2\]

OpenStudy (anonymous):

sorry for the late response, but that was the answer in my textbook to a problem. I was thinking -ln2 / 2^x but the text says -1/2^x ln2 which is why I wanted to double check my algebra here

OpenStudy (anonymous):

the problem was to find the integral of 2^-x

OpenStudy (anonymous):

Since it's been so long since I took calculus. Why don't post the intergal of 2^-x as a new question. Maybe there is something you are not seeing.

OpenStudy (anonymous):

i just figured it out by looking at it again roflmao

OpenStudy (anonymous):

it was the calculus, not the algebra, damn!